Standard Costing

Botella Company produces plastic bottles. The unit for costing purposes is a case of 18 bottles. The following standards for producing one case of bottles have been established:

Direct materials (4 lbs. @ $0.95) $ 3.80
Direct labor (1.25 hours @ $15.00) 18.75
Standard prime cost 22.55

During December, 78,000 pounds of material were purchased and used in production. There were 15,000 cases produced, with the following actual prime costs:

Direct materials $ 74,000
Direct labor $315,000 (for 22,500 hrs.)

Required:

1. Compute the materials variances. If there is no variance, enter "0" in the amount box and select "No variance" from the dropdown menu.

Materials price variance $
Materials usage variance $
Overall materials variance $

2. Compute the labor variances. If there is no variance, enter "0" in the amount box and select "No variance" from the dropdown menu.

Labor rate variance $
Labor efficiency variance $
Overall labor variance $
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (0.95-74,000/78,000)*78000

100 F

Materials usage variance = (SQ-AQ) *SP

= (15,000*4 - 78,000)*0.95

17,100 U
Overall materials variance 17,000 U

.

Labor Rate Variance = (SR-AR) *AH

= (15-315000/22500)*22500

22500 F

Labor efficiency variance = (SH-AH) *SR

= (15,000*1.25 - 22500)*15

56,250 U
Overall variance 33,750 U

Comment if you face any issues